# Density estimation via mixture discrepancy and moments

### v1

@@ 4.4 $d$-D Mixtures of Gaussians
where $\boldsymbol{\alpha}=(0.4,0.3,0.2,0.1)$, $\boldsymbol{\mu}_{1}=0.3\mathbbm{1},\boldsymbol{\mu}_{2}=0.4\mathbbm{1},\boldsymbol{\mu}_{3}=0.5\mathbbm{1},\boldsymbol{\mu}_{4}=0.6\mathbbm{1}$, $\boldsymbol{\Sigma}_{1}=0.01\boldsymbol{I},\boldsymbol{\Sigma}_{2}=0.02\boldsymbol{I},\boldsymbol{\Sigma}_{3}=0.01\boldsymbol{I},\boldsymbol{\Sigma}_{4}=0.02\boldsymbol{I}$. $\mathbbm{1}$ is a $1\times d$ all-one vector, and $\boldsymbol{I}$ is a $d\times d$ identity matrix. Table 4 displays the numerical errors and computational costs of the three methods under different sample sizes $N$ and dimensions $d$. The numerical results at different dimensions are still similar to those of the previous 2-D examples. The $L^{2}$ relative errors $\mathcal{E}_{2}[p]$ of the three density estimations are generally similar. However, in terms of running time, DSP-mix and MSP are significantly faster and can generally achieve a tenfold speedup. When the number of particles is relatively large (e.g. $N=1\times 10^{6}$), MSP exhibits greater speed advantages over DSP-mix. The underlying reason lies in the fact that the computational complexity of MSP scales linearly with the number of particles, in contrast to that of DSP-mix which is quadratic.
Table 4: $d$-D Mixtures of Gaussians: The $L^{2}$ relative error $\mathcal{E}_{2}[p]$ and running time in seconds (in parentheses) of the three methods under different sample sizes $N$ and dimensions $d$.
$N=1\times 10^{4}$ $N=1\times 10^{5}$ $N=1\times 10^{6}$
$d$ DSP DSP-mix MSP DSP DSP-mix MSP DSP DSP-mix MSP
2 0.2340 0.2106 0.2489 0.0991 0.0723 0.0771 0.0415 0.0231 0.0211
(0.7537) (0.0558) (0.0133) (2.6471) (0.2065) (0.1799) (13.30) (2.4144) (2.1306)
3 0.2966 0.2237 0.5455 0.1341 0.1268 0.1733 0.0773 0.0663 0.0615
(2.1429) (0.0189) (0.0187) (7.3443) (0.3405) (0.2131) (57.00) (6.9713) (2.823)
4 0.3748 0.3108 0.7009 0.2211 0.2374 0.2715 0.1672 0.1582 0.1292
(3.0253) (0.0208) (0.0209) (10.1659) (0.3949) (0.2517) (67.59) (9.184) (3.4781)
5 0.4864 0.4700 0.7270 0.3530 0.4095 0.3734 0.2340 0.3006 0.2220
(3.2468) (0.0246) (0.024) (12.8835) (0.4635) (0.2914) (82.18) (10.94) (4.3056)
6 0.6148 0.7065 0.7792 0.4870 0.5746 0.4623 0.2966 0.4578 0.3227
(3.8357) (0.0291) (0.0282) (15.8595) (0.5347) (0.3243) (99.17) (12.78) (5.0615)
4.5 $d$-D Mixtures of Betas
Finally we examine the $d$-D mixtures of betas,
Table 5 displays the numerical errors and computational costs of the three methods for $d$-D mixtures of betas case. The errors of the three methods are comparable, all decreasing with the increase of the sample number $N$ and increasing with the rise of the dimension $d$. When $N\geq 1\times 10^{5}$, DSP-mix is approximately ten times faster than DSP, while MSP attains an acceleration approximately a hundred times. Figure 6 illustrates the influence of the parameters on the DSP-mix method and the MSP method. When $\theta,\epsilon_{1},\epsilon_{2}$ and $\epsilon_{3}$ decrease within a certain range, we can obtain finer partitioning and more accurate density estimation results. However, when these parameters are too small, the problem of overfitting will occur, and the error will instead increase. For the number of samples $N$, its increase will definitely make the solution more accurate.
Table 5: $d$-D Mixtures of Betas: The $L^{2}$ relative error $\mathcal{E}_{2}[p]$ and running time in seconds (in parentheses) of the three methods under different sample sizes $N$ and dimensions $d$.
$N=1\times 10^{4}$ $N=1\times 10^{5}$ $N=1\times 10^{6}$
$d$ DSP DSP-mix MSP DSP DSP-mix MSP DSP DSP-mix MSP
2 0.2232 0.1954 0.2875 0.1077 0.0841 0.0846 0.0457 0.0274 0.0258
(0.9318) (0.0165) (0.0080) (3.6159) (0.2168) (0.0494) (17.1128) (2.2811) (0.5550)
3 0.2790 0.2319 0.5789 0.1312 0.1241 0.1851 0.0818 0.0690 0.0634
(2.4086) (0.0191) (0.0173) (8.7814) (0.3537) (0.0900) (61.5023) (7.0977) (0.9356)
4 0.3720 0.3224 0.7096 0.2318 0.2439 0.2908 0.1825 0.1694 0.1344
(3.2390) (0.0223) (0.0238) (11.3189) (0.4173) (0.1315) (76.6812) (8.9815) (1.2242)
5 0.4646 0.4561 0.7782 0.3326 0.3838 0.3713 0.2926 0.2964 0.2209
(3.6802) (0.0250) (0.0314) (14.6606) (0.4746) (0.1701) (91.7739) (10.3924) (1.5386)
6 0.5597 0.5910 0.7828 0.4555 0.5355 0.4423 0.4122 0.4311 0.3154
(4.2078) (0.0280) (0.0376) (16.0885) (0.5236) (0.2034) (104.2499) (11.9599) (1.8635)
(a) The influence of $\theta$ on the error of DSP-mix.
(b) The influence of $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon$ on the error of MSP.
[…]

### v2

@@ 4.2 $d$-D Gaussian mixtures
where $\boldsymbol{\alpha}=(0.4,0.3,0.2,0.1)$, $\boldsymbol{\mu}_{1}=0.3\mathbbm{1},\boldsymbol{\mu}_{2}=0.4\mathbbm{1},\boldsymbol{\mu}_{3}=0.5\mathbbm{1},\boldsymbol{\mu}_{4}=0.6\mathbbm{1}$, $\boldsymbol{\Sigma}_{1}=0.01\boldsymbol{I},\boldsymbol{\Sigma}_{2}=0.02\boldsymbol{I},\boldsymbol{\Sigma}_{3}=0.01\boldsymbol{I},\boldsymbol{\Sigma}_{4}=0.02\boldsymbol{I}$. Here $\mathbbm{1}$ is a $1\times d$ all-one vector, and $\boldsymbol{I}$ is a $d\times d$ identity matrix.
(a) KL divergence ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 8: 22-D Gaussian mixtures: The KL divergence and Hellinger distance under different NN and $\theta$.
(a) True density and samples.
(b) Adaptive partition (left: DSP, middle: DSP-mix, right: MSP).
(c) Density estimation (left: DSP, middle: DSP-mix, right: MSP).
Figure 9: 2-D Gaussian mixtures: Adaptive partitions and density estimators produced by DSP-mix, MSP and DSP with $N=1\times 10^{4}$ and $\theta=0.2$.
Figure 9 presents the numerical solutions of DSP, DSP-mix and MSP. All of them possess the capability to capture the distribution structure. They feature fine grids at positions where the probability density changes significantly and sparse grids in unimportant regions.
From Figure 8, the V-shape curves with respect to $\theta$ are still observed in both error metrics. Therefore, it is suggested to avoid using too small parameter $\theta$ unless the sample size is large enough.
For high-dimensional tests under $d=15,20,30$, Figure 10 presents the convergence of error metrics with respect to sample size $N=10^{4},10^{5},10^{6},10^{7},10^{8}$, with the raw data collected in Tables 5 and 7. Figure 11 plots the partition level under different $d$, NN, with $\theta$ fixed to be $0.002$, with the raw data collected in Tables 6 and 8.
For the rapidly decreasing Gaussian distribution, the performance of DSP, DSP-mix and MSP is very similar to those for the localized Beta mixtures under different dimensions. It is still observed that the accuracy of all density estimators can be improved by increasing either sample size $N$ or increasing the partition level $L$. In terms of running time, DSP-mix and MSP are significantly faster and can generally achieve a tenfold or even twenty-fold speedup for large sample size ($N\geq 10^{7}$). Moreover, the trend $L\sim\theta^{-1}$ is still observed in all cases with only exception $d=2$.
(a) KL divergence (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level (left: DSP, middle: DSP-mix, right: MSP).
Figure 10: $d$-D Gaussian mixtures: Convergence of DSP, DSP-mix and MSP with respect to NN for $d=15,20,30$.
(a) Partition level ($N=10^{4}$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Partition level ($N=10^{5}$) (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level ($N=10^{6}$) (left: DSP, middle: DSP-mix, right: MSP).
(d) Partition level ($N=10^{7}$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 11: $d$-D Gaussian mixture: The partition level $L$ under different $\theta$, $N$ and $d$.
Table 5: $d$-D Gaussian mixtures: The KL divergence and Hellinger distance under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
2 DSP -0.0008 0.0005 0.0014 0.0007 0.0020 0.0008 0.0030 0.0010 0.0042 0.0012
DSP-mix 0.0002 0.0009 0.0020 0.0009 0.0027 0.0010 0.0051 0.0016 0.0095 0.0028
MSP -0.0009 0.0005 0.0012 0.0007 0.0018 0.0007 0.0025 0.0009 0.0034 0.0011
(maxSD) 0.0001 0.0000 0.0001 0.0000 0.0001 0.0000 0.0002 0.0001 0.0003 0.0001
4 DSP 0.9405 0.0551 0.0046 0.0074 0.0187 0.0100 0.0356 0.0128 0.0510 0.0162
DSP-mix -0.0621 0.0173 0.0125 0.0092 0.0338 0.0128 0.0597 0.0195 0.0852 0.0250
MSP 0.8370 0.0798 -0.0118 0.0076 0.0062 0.0076 0.0164 0.0090 0.0262 0.0107
(maxSD) 3.4708 0.0026 0.0189 0.0026 0.0284 0.0026 0.0294 0.0026 0.0292 0.0027
6 DSP -0.3613 0.0969 0.0830 0.0284 0.1336 0.0368 0.1671 0.0437 0.2146 0.0547
DSP-mix -0.0253 0.0304 0.1213 0.0348 0.2003 0.0510 0.2791 0.0669 0.3264 0.0765
MSP -0.4532 0.1162 -0.0007 0.0271 0.0608 0.0273 0.0960 0.0308 0.1261 0.0355
(maxSD) 0.0003 0.0004 0.0008 0.0004 0.0023 0.0006 0.0029 0.0008 0.0034 0.0008
8 DSP -0.4183 0.1383 0.2277 0.0676 0.3186 0.0817 0.3812 0.0935 0.4916 0.1177
DSP-mix 0.0799 0.0626 0.3278 0.0832 0.5159 0.1182 0.6375 0.1377 0.7058 0.1510
MSP -0.4558 0.1455 0.0833 0.0623 0.2030 0.0667 0.2674 0.0734 0.3281 0.0831
(maxSD) 0.0011 0.0006 0.0025 0.0008 0.0036 0.0006 0.0047 0.0010 0.0058 0.0013
10 DSP -0.3841 0.1836 0.4510 0.1210 0.5888 0.1395 0.7002 0.1602 0.9045 0.1979
DSP-mix 0.2744 0.1115 0.6628 0.1523 0.9963 0.2029 1.1433 0.2224 1.2483 0.2418
MSP -0.3919 0.1848 0.2745 0.1113 0.4514 0.1217 0.5395 0.1313 0.6628 0.1523
(maxSD) 0.0021 0.0006 0.0036 0.0008 0.0061 0.0014 0.0122 0.0015 0.0118 0.0017
15 DSP 0.0593 0.3195 1.4166 0.2948 1.7380 0.3296 2.1080 0.3816 2.5918 0.4280
DSP-mix 1.2203 0.2818 2.2709 0.3894 2.8796 0.4337 3.1056 0.4576 3.5545 0.5211
MSP 0.0593 0.3198 1.2203 0.2817 1.5263 0.3001 1.7615 0.3296 2.2709 0.3894
(maxSD) 0.0053 0.0015 0.0081 0.0014 0.0205 0.0027 0.0264 0.0022 0.0295 0.0028
20 DSP 1.0081 0.4844 3.0613 0.4864 3.7281 0.5460 4.4446 0.5987 5.1962 0.6306
DSP-mix 2.8806 0.4746 4.9634 0.6092 5.6702 0.6371 6.1985 0.6827 7.3848 0.7647
MSP 1.0081 0.4843 2.8806 0.4748 3.3849 0.5061 4.0691 0.5638 4.9634 0.6094
(maxSD) 0.0200 0.0017 0.0311 0.0024 0.0483 0.0029 0.0591 0.0027 0.0668 0.0028
30 DSP 4.8871 0.7676 8.5323 0.7894 10.1693 0.8359 11.5031 0.8534 12.6868 0.8754
DSP-mix 8.3667 0.7800 12.4234 0.8571 13.8661 0.8963 15.4386 0.9284 17.0640 0.9359
MSP 4.8871 0.7679 8.3667 0.7805 9.9671 0.8262 11.3469 0.8446 12.4234 0.8570
(maxSD) 0.0320 0.0012 0.0418 0.0014 0.0426 0.0017 0.0555 0.0015 0.0612 0.0013
Table 6: $d$-D Gaussian mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
2 DSP 6603 7.88 3372 13.35 2302 18.66 1708 27.79 1239 45.49
DSP-mix 8353 6.92 3315 7.38 2260 9.03 1672 11.67 1200 20.86
MSP 7001 7.28 3608 8.83 2397 9.03 1788 8.99 1309 9.18
4 DSP 853559 882.76 27793 269.96 11255 270.44 5823 296.21 3155 323.50
DSP-mix 255266 21.33 25823 25.91 10390 47.30 5178 81.74 2649 149.73
MSP 1296208 32.90 62370 17.15 24249 14.46 14210 14.08 8777 13.24
6 DSP 1372339 1109.13 40997 426.41 16269 385.83 8894 398.37 4900 429.82
DSP-mix 271883 25.15 27674 36.71 11259 67.65 5568 118.32 2902 215.97
MSP 1732952 41.91 193079 26.18 78787 22.51 42202 20.70 23052 19.07
8 DSP 1711429 937.34 63630 613.12 25090 523.57 13821 512.66 7422 538.91
DSP-mix 287074 28.99 30227 47.96 11765 88.58 5933 152.11 3350 272.48
MSP 1850228 49.01 276904 34.20 110642 30.05 56644 27.22 29844 24.78
10 DSP 1927372 798.32 89458 774.72 35850 633.31 20018 606.81 10431 634.64
DSP-mix 301812 36.70 32765 58.30 12257 107.54 6529 181.35 3899 326.22
MSP 1955139 56.54 301284 38.56 119087 34.70 62357 31.48 32749 29.19
15 DSP 2268841 663.98 166721 988.71 72700 825.14 39392 851.34 19019 800.97
DSP-mix 336290 55.35 38372 85.67 14657 149.58 9097 243.25 5480 464.71
MSP 2268874 78.76 336290 54.82 143598 49.61 80693 45.85 38372 41.86
20 DSP 2721235 639.02 269423 1184.09 120788 1148.42 61092 1131.69 28893 914.53
DSP-mix 395153 77.47 42435 111.01 19350 185.18 12237 309.70 6016 613.64
MSP 2721235 105.96 395153 73.37 189409 66.72 96443 60.69 42435 55.44
30 DSP 3754332 734.10 526282 1923.58 212475 1829.85 99414 1466.28 50224 1479.35
DSP-mix 601142 115.55 59909 159.59 28074 265.94 13985 458.06 6604 843.57
MSP 3754333 167.07 601142 113.35 247671 98.94 114555 89.74 59909 85.51
Table 7: $d$-D Gaussian mixtures: The KL divergence and Hellinger distance under different dimensions $d$ and sample size $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
15 DSP 2.3652 0.6851 1.3628 0.5327 0.5726 0.4028 0.0593 0.3195 0.4881 0.2012
DSP-mix 2.3652 0.6853 1.3628 0.5332 1.1873 0.3716 1.2203 0.2818 0.9694 0.2203
MSP 2.3652 0.6848 1.3628 0.5331 0.5726 0.4033 0.0593 0.3198 0.2821 0.2201
20 DSP 4.9939 0.8507 3.5812 0.7394 2.0609 0.5949 1.0081 0.4844 1.2630 0.3555
DSP-mix 4.9939 0.8499 3.5812 0.7393 2.9840 0.5818 2.8806 0.4746 2.3519 0.4071
MSP 4.9939 0.8516 3.5812 0.7395 2.0609 0.5945 1.0081 0.4843 1.1115 0.3652
30 DSP 11.3623 0.9714 9.6601 0.9375 7.2038 0.8627 4.8871 0.7676 4.3272 0.6549
DSP-mix 11.3623 0.9717 9.6601 0.9374 9.1441 0.8652 8.3667 0.7800 7.3474 0.7180
MSP 11.3623 0.9713 9.6601 0.9376 7.2038 0.8626 4.8871 0.7679 4.2576 0.6557
Table 8: $d$-D Gaussian mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
15 DSP 3383 0.03 28693 0.37 246147 5.42 2268841 663.98 5838142 30802.07
DSP-mix 3383 0.03 28693 0.44 126396 5.11 336290 55.35 1042368 803.32
MSP 3383 0.03 28693 0.31 246147 5.49 2268874 78.76 10250539 752.33
20 DSP 3775 0.04 34852 0.57 305394 7.96 2721235 639.02 8503962 34682.44
DSP-mix 3775 0.04 34852 0.64 162165 6.76 395153 77.47 1170087 1076.32
MSP 3775 0.04 34852 0.48 305394 7.84 2721235 105.96 11362133 882.67
30 DSP 4433 0.07 46950 0.86 441887 13.93 3754332 734.10 14664912 42827.37
DSP-mix 4433 0.07 46950 1.04 217169 11.25 601142 115.55 1485038 1677.59
MSP 4433 0.07 46950 0.98 441887 13.49 3754333 167.07 15638938 1439.79
4.3 $d$-D Cauchy Mixtures
Finally, we consider a much more challenging problem: Learning the density of the multidimensional Cauchy mixtures [9]
$\bm{x}_{i}\sim\sum_{j=1}^{q}p_{j}C(\bm{x},\bm{m}_{j},\bm{u}_{j}),\quad C(\bm{x},\bm{m}_{j},\bm{u}_{j})=\prod_{k=1}^{d}\frac{u_{jk}}{\pi\left[u_{jk}^{2}+(x_{k}-m_{jk})^{2}\right]}.$ (35)
All the parameters are set to be: $q=4$, $p_{1}=p_{2}=p_{3}=p_{4}=0.25$, $u_{jk}=0.32$ for all $j,k$ and the centres $m_{1k}=-3$, $m_{2k}=-1$, $m_{3k}=1$, $m_{4k}=3$ for $k=1,\dots,d$.
(a) KL divergence ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 12: 22-D Cauchy mixtures: The KL divergence and Hellinger distance under different NN and $\theta$.
(a) True density and samples.
(b) Heavy tail of Cauchy mixtures.
(c) Adaptive partition (left: DSP, middle: DSP-mix, right: MSP).
(d) Density estimation (left: DSP, middle: DSP-mix, right: MSP).
Figure 13: 2-D Cauchy mixtures: Adaptive partitions and density estimators produced by DSP-mix, MSP and DSP with $N=1\times 10^{4}$ and $\theta=0.2$.
One can sample the Cauchy distribution by using its cumulative function
$F\left(\frac{x}{u_{j}}\right)=\frac{1}{\pi}\left[\arctan\left(\frac{x}{u_{j}}-\frac{m_{j}}{u_{j}}\right)+\frac{\pi}{2}\right].$ (36)
As shown in Figure 13, the Cauchy distribution has a very heavy tail, although the typical set lies in a bounded domain $15,15]^{d}$. In fact, the Cauchy distribution has infinite mean and variance, so that direct Monte Carlo sampling may fail to converge. Therefore, we only try to reconstruct the density in the typical set $\Omega=[-15,15]^{d}$ and discard the particles outside $\Omega$.
Indeed, the heavy tail poses a significant challenge for the tree-based density estimation. Nevertheless, DSP, DSP-mix and MSP are able to capture the four-peak structure in the typical set, as clearly visualized in Figure 13. Figure 12 shows that the behavior of the KL divergence is a bit strange compared with Figures 4 and 8 for 2-D Beta mixtures and 2-D Gaussian mixtures. It may be also related to the heavy tail and we guess that, when the domain is enlarged, the performance of tree-based density estimator should be more sensitive to the partition level. As $\theta$ increases, the decision tree stops at a shallow level, producing coarser but more stable regions. This coarsening effectively suppresses the influence of isolated Cauchy outliers, causing the KL divergence to be smaller. That is, the KL divergence no longer serves as a satisfactory criterion for such heavy tail distribution as also pointed out in [32]. By contrast, the Hellinger distance always displays a similar behavior for the three mixtures.
For $d=15,20,30$, Figure 14 presents the convergence of error metrics with respect to sample size $N=10^{4},10^{5},10^{6},10^{7},10^{8}$. The raw data is collected in Tables 9 and 7. Figure 15 plots the partition level $L$ under different $d$, NN, with $\theta$ fixed to be $0.002$. The raw data is collected in Tables 10 and 12.
  1. (1)
    Overall, the errors diminish as either sample size $N$ and the partition level $L$ increase. The accuracy of MSP is comparable to DSP, whereas DSP-mix is less accurate as dimension $d$ increases.
  2. (2)
    Notably, for higher-dimensional cases with larger-sample, e.g., $d=30,N=10^{7}$, the KL divergence behaves smoothly across all $\theta$, confirming that the estimator can work well provided sufficient sample coverage of the heavy-tailed support.
  3. (3)
    The trend $L\sim\theta^{-1}$ is still observed in Figure 15. This gives us a simple way to estimate the partition level by first testing on a small dataset and using extrapolation, which also helps control the computational complexity.
(a) KL divergence (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level (left: DSP, middle: DSP-mix, right: MSP).
Figure 14: $d$-D Cauchy mixtures: Convergence of DSP, DSP-mix and MSP with respect to NN for $d=15,20,30$.
(a) Partition level ($N=10^{4}$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Partition level ($N=10^{5}$) (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level ($N=10^{6}$) (left: DSP, middle: DSP-mix, right: MSP).
(d) Partition level ($N=10^{7}$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 15: $d$-D Cauchy mixtures: The partition level $L$ under different $\theta$, $N$ and $d$.
Table 9: $d$-D Cauchy mixtures: The KL divergence and Hellinger distance under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
2 DSP 0.4771 0.1892 0.3830 0.1789 0.3836 0.1789 0.3765 0.1795 0.3425 0.1806
DSP-mix 0.4229 0.1844 0.3832 0.1789 0.3841 0.1790 0.3777 0.1798 0.3454 0.1814
MSP 0.4414 0.1962 0.3825 0.1790 0.3832 0.1788 0.3755 0.1792 0.3407 0.1801
(maxSD) 0.1507 0.0002 0.0325 0.0001 0.0190 0.0001 0.0121 0.0001 0.0182 0.0001
4 DSP 0.1406 0.2198 0.1003 0.1922 0.1190 0.2019 0.0520 0.2128 0.1000 0.2255
DSP-mix 0.1276 0.1888 0.1334 0.2015 0.0730 0.2164 0.1408 0.2350 0.0086 0.2573
MSP 0.1231 0.2460 0.0724 0.1866 0.1129 0.1887 0.1199 0.1936 0.1312 0.2018
(maxSD) 0.2531 0.0001 0.0735 0.0001 0.0539 0.0001 0.0383 0.0001 0.0864 0.0003
6 DSP -0.1851 0.2715 -0.0588 0.2648 -0.1155 0.2925 -0.0748 0.3189 -0.0257 0.3481
DSP-mix -0.0882 0.2456 0.0001 0.3083 0.0902 0.3492 0.1615 0.3862 0.2793 0.4290
MSP -0.2477 0.2841 -0.1375 0.2455 -0.1319 0.2619 -0.0682 0.2813 -0.0062 0.3083
(maxSD) 0.1513 0.0001 0.1260 0.0003 0.1220 0.0004 0.1261 0.0005 0.0551 0.0007
8 DSP -0.6951 0.3909 -0.2798 0.3852 -0.1054 0.4298 0.0831 0.4677 0.1232 0.5088
DSP-mix -0.3676 0.3578 0.1077 0.4634 0.2991 0.5203 0.4819 0.5630 0.8625 0.6147
MSP -0.7145 0.3640 -0.3676 0.3578 -0.1418 0.3914 -0.0605 0.4254 0.1077 0.4634
(maxSD) 0.3285 0.0888 0.3285 0.0005 0.1064 0.0005 0.1113 0.0006 0.0973 0.0007
10 DSP -0.9253 0.4730 -0.0914 0.5246 0.1218 0.5767 0.4481 0.6182 0.5655 0.6593
DSP-mix -0.1747 0.4957 0.5718 0.6209 0.8718 0.6745 1.2397 0.7188 1.5495 0.7573
MSP -0.9281 0.4732 -0.1747 0.4958 0.0123 0.5399 0.2138 0.5793 0.5718 0.6209
(maxSD) 0.2714 0.0002 0.1065 0.0003 0.0482 0.0005 0.0852 0.0005 0.0401 0.0003
15 DSP -0.6121 0.7473 1.0220 0.8125 1.7181 0.8518 2.2469 0.8775 2.8725 0.9002
DSP-mix 0.9721 0.7955 2.6561 0.8834 3.4443 0.9143 4.0266 0.9337 4.5270 0.9455
MSP -0.6122 0.7473 0.9721 0.7955 1.6077 0.8333 2.1442 0.8597 2.6561 0.8834
(maxSD) 0.1532 0.0003 0.0943 0.0004 0.0908 0.0004 0.0768 0.0004 0.0234 0.0003
20 DSP 0.7989 0.9095 3.5479 0.9432 4.6367 0.9587 5.3676 0.9682 6.2838 0.9773
DSP-mix 3.5414 0.9382 6.1021 0.9721 7.1799 0.9818 7.9012 0.9856 8.4159 0.9892
MSP 0.7989 0.9095 3.5414 0.9382 4.5788 0.9533 5.2970 0.9629 6.1021 0.9721
(maxSD) 0.1970 0.0002 0.0908 0.0002 0.0681 0.0002 0.0884 0.0002 0.0258 0.0002
30 DSP 6.7636 0.9927 11.2454 0.9963 12.8170 0.9978 14.0692 0.9986 15.1920 0.9990
DSP-mix 11.2116 0.9960 15.1105 0.9988 16.4078 0.9992 17.1863 0.9996 18.0573 0.9998
MSP 6.7636 0.9927 11.2116 0.9960 12.7554 0.9975 13.9290 0.9984 15.1105 0.9988
(maxSD) 0.2761 0.0001 0.1189 0.0001 0.1874 0.0000 0.1519 0.0000 0.1898 0.0000
Table 10: $d$-D Cauchy mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
2 DSP 173841 151.72 16211 87.81 6639 105.19 3332 126.31 1691 146.15
DSP-mix 112188 8.84 16154 9.98 6583 16.92 3274 28.84 1625 52.87
MSP 280402 8.83 18032 6.07 6925 5.63 3698 5.25 2059 5.12
4 DSP 648039 687.78 27136 210.24 11910 192.20 6492 196.96 3659 205.85
DSP-mix 175907 14.01 17613 17.35 7307 29.68 3608 51.81 1890 94.56
MSP 1043302 17.84 115907 9.39 52032 9.07 28907 8.62 15741 8.29
6 DSP 1006568 678.35 53796 309.37 23027 257.94 12423 244.18 6906 245.08
DSP-mix 191134 18.10 19539 24.00 8038 42.36 4164 73.33 2158 137.66
MSP 1236975 22.90 188354 13.62 76132 12.40 38998 11.21 19520 10.63
8 DSP 1325701 553.24 85691 375.92 35788 290.85 18973 274.74 10181 272.09
DSP-mix 209324 23.47 21821 31.33 8940 54.64 4658 95.43 2379 177.23
MSP 1390910 29.26 209290 17.38 84843 16.10 43202 14.63 21821 13.77
10 DSP 1546736 461.04 119372 408.73 49117 319.62 25449 301.88 13155 296.88
DSP-mix 229815 28.18 23945 37.43 9809 66.84 5086 116.65 2651 210.02
MSP 1555407 33.94 229815 21.78 93768 18.77 47998 17.08 23945 16.16
15 DSP 1977162 340.83 204560 511.87 80332 421.72 39741 405.71 19590 395.28
DSP-mix 286280 38.17 28866 54.95 11644 97.07 5792 164.47 3199 290.12
MSP 1977232 51.67 286280 33.44 116620 28.26 58439 26.55 28866 24.54
20 DSP 2399337 328.32 279421 636.23 106726 576.62 51803 560.24 24388 534.52
DSP-mix 334965 53.23 32580 71.41 12502 126.16 6370 212.48 3633 363.47
MSP 2399363 71.01 334964 44.12 135152 37.84 67249 36.07 32580 32.42
30 DSP 3125241 386.49 389512 890.28 148772 890.74 68230 867.85 30332 839.55
DSP-mix 407694 81.72 35503 104.93 13466 182.92 7247 298.40 4264 513.47
MSP 3125248 109.90 407670 66.05 164124 59.87 77742 52.82 35503 48.01
Table 11: $d$-D Cauchy mixtures: The KL divergence and Hellinger distance under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
15 DSP 3.4022 0.9513 1.4427 0.9019 0.1989 0.8318 -0.6121 0.7473 1.0773 0.6183
DSP-mix 3.4022 0.9513 1.4427 0.9018 1.0771 0.8454 0.9721 0.7955 2.7906 0.7060
MSP 3.4022 0.9514 1.4427 0.9018 0.1739 0.8319 -0.6122 0.7473 2.7906 0.7062
20 DSP 7.1493 0.9904 4.3705 0.9767 2.4521 0.9500 0.7989 0.9095 2.7486 0.8506
DSP-mix 7.1493 0.9904 4.3705 0.9767 3.7805 0.9567 3.5414 0.9382 5.2577 0.9047
MSP 7.1493 0.9904 4.3705 0.9767 2.4521 0.9500 0.7989 0.9095 5.2577 0.9047
30 DSP 15.9541 0.9997 12.5253 0.9990 9.2726 0.9970 6.7636 0.9927 6.3239 0.9863
DSP-mix 15.9541 0.9997 12.5253 0.9990 11.4016 0.9977 11.2116 0.9960 10.3196 0.9940
MSP 15.9541 0.9997 12.5253 0.9990 9.2726 0.9970 6.7636 0.9927 10.3196 0.9940
Table 12: $d$-D Cauchy mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
15 DSP 2843 0.03 24527 0.25 222954 4.33 1977162 340.83 7840212 19084.37
DSP-mix 2843 0.03 24527 0.26 106704 3.53 286280 38.17 987999 596.05
MSP 2843 0.02 24527 0.24 222960 4.46 1977232 51.67 987999 570.70
20 DSP 3431 0.04 29100 0.37 272278 6.39 2399337 328.32 10550658 20532.97
DSP-mix 3431 0.04 29100 0.38 127445 4.97 334965 53.23 1086785 716.39
MSP 3431 0.04 29100 0.35 272278 6.36 2399363 71.01 1086785 732.24
[…]